OpenStudy (anonymous):

Tammy can clean half of an area in an hour. Melissa can clean one-third of the area in an hour. How much of the area can they clean in 45 minutes?

OpenStudy (jojokiw3):

Rate of Tammy (Time of Tammy) + Rate of Melissa (Time of Melissa) = Area Cleaned. (Time of Tammy) = (Time of Melissa) = 45. Rate of Tammy = \[\frac{ \frac{ 1 }{ 2 }area }{ 60 min }\] Rate of Melissa = \[\frac{ \frac{ 1 }{ 3 }area }{ 60 min}\]

OpenStudy (anonymous):

And then what?

OpenStudy (jojokiw3):

So you plug things in and do the first equation I mentioned.

OpenStudy (jojokiw3):

Remember to cancel things out too, including units. :)

OpenStudy (anonymous):

I don't get it though. Usually, I'd be able to, but right now they just look like puzzle pieces that don't fit.

OpenStudy (jojokiw3):

hmm. Okay. Lemme see if I can rephrase. :P

OpenStudy (anonymous):

Thanks, that might help. :|

OpenStudy (jojokiw3):

Okay, so Rate = how fast you can do something. How fast can Tammy clean half of an area?

OpenStudy (anonymous):

In 30 minutes

OpenStudy (jojokiw3):

So \[R = \frac{ area }{ time }\] In this case.

OpenStudy (jojokiw3):

Thereby, in numbers, what is Tammy's Rate exactly?

OpenStudy (anonymous):

Uhm, 30 minutes?

OpenStudy (jojokiw3):

Ahh sorry my bad. Lemme say that again. Rate = \[\frac{ Area }{ Time }\] Tammy cleans half an area in 1 hour. So divide that and you get the rate.

OpenStudy (jojokiw3):

Rate of Tammy, at least.

OpenStudy (anonymous):

1/120

OpenStudy (jojokiw3):

1 in 2 hours or 120 minutes. Both correct. Let's switch it back to hours though, just for simpleness. So \[\frac{ 1area }{ 2hrs }\]

OpenStudy (jojokiw3):

Do the same thing for Melissa. So \[\frac{ \frac{ 1 }{ 3 }area }{ 1hr }\]

OpenStudy (anonymous):

20

OpenStudy (jojokiw3):

Wait. 20 = \[\frac{ \frac{ 1 }{ 3 }area }{ 1 hr}?\]

OpenStudy (anonymous):

No. I just found 1/3 of 60 though, which is 20.

OpenStudy (jojokiw3):

:P Forget the minutes, again. Simply just do the division with hours.

OpenStudy (anonymous):

Oh, :P So it'd be 1 area in 3 hours?

OpenStudy (jojokiw3):

Yep!

OpenStudy (jojokiw3):

now 45 minutes is what fraction of an hour?

OpenStudy (anonymous):

3/4

OpenStudy (jojokiw3):

Now multiply both the rate of tammy and the rate of melissa with that, then add the two resultants and you have your answer.

OpenStudy (jojokiw3):

So 1/2(3/4) + 1/3(3/4) = answer. Remember to cancel units out.

OpenStudy (anonymous):

1/2(3/4) + 1/3(3/4) = 9/12 or 3/4

OpenStudy (jojokiw3):

wait what? 1/2(3/4) = 3/8. whereas 1/3(3/4) = 3/12. Get the LCM of those and add.

OpenStudy (anonymous):

Oh, I multiplied wrong. :P Okay so it'd be 9/24 + 6/24 = 15/24 ??? I feel like this makes less sense now.

OpenStudy (jojokiw3):

Yep, now simply simplify that into 5/8 and you're right.

OpenStudy (jojokiw3):

So memorize those equations and plug things in and you should be good on these questions for the rest of your life.

OpenStudy (anonymous):

Oh, I didn't simplify. :P

OpenStudy (anonymous):

Okay wait, so they cleaned 5/8 of the area in 45 minutes?

OpenStudy (jojokiw3):

Yep. :P

OpenStudy (anonymous):

Okay, now it makes sense, thanks for all your help. :)

OpenStudy (jojokiw3):

No problem. :)